# SplitVAEs: Decentralized scenario generation from siloed data for stochastic optimization problems

### v1

@@ Abstract
Stochastic optimization problems in large-scale multi-stakeholder networked systems (e.g., power grids and supply chains) rely on data-driven scenarios to encapsulate complex spatiotemporal interdependencies. However, centralized aggregation of stakeholder data is challenging due to the existence of data silos resulting from computational and logistical bottlenecks. In this paper, we present SplitVAEs, a decentralized scenario generation framework that leverages variational autoencoders to generate high-quality scenarios without moving stakeholder data. With the help of experiments on distributed memory systems, we demonstrate the broad applicability of SplitVAEs in a variety of domain areas that are dominated by a large number of stakeholders. Our experiments indicate that SplitVAEs can learn spatial and temporal interdependencies in large-scale networks to generate scenarios that match the joint historical distribution of stakeholder data in a decentralized manner. Our experiments show that SplitVAEs deliver robust performance compared to centralized, state-of-the-art benchmark methods while significantly reducing data transmission costs, leading to a scalable, privacy-enhancing alternative to scenario generation.
@@ I Introduction
Stochastic optimization (SO) problems are essential components in large-scale, data-driven planning and decision frameworks for multi-stakeholder-led networked systems such as energy [1, 2], supply chain [3], spare parts management [4], healthcare [5] and logistics [6]. Naturally, good quality solutions to SO problems require high-fidelity data-driven scenarios that encapsulate spatially and temporally correlated interdependencies from myriad sources. As a result, generating high-fidelity scenarios depends on the availability of high-quality datasets that capture these interdependencies from heterogeneous data sources. However, conventional scenario generation methods necessitate centralized data aggregation to capture interdependencies to drive stochastic planning decisions. Centralized aggregation schemes for scenario generation face several computational, and logistical challenges [7, 8] due to the presence of data silos across multiple stakeholders of a networked system. Therefore, in this study, we present SplitVAEs, a novel decentralized scenario generation framework that utilizes variational autoencoders (VAE) to eliminate centralized data aggregation. SplitVAEs are compatible with pre-existing data silos, ensure scalable generation of high-fidelity scenarios and significantly reduce data transfer costs.
Conventionally the process of solving an SO problem begins with the acquisition of local spatiotemporally correlated datasets from individual stakeholders. The mechanism of data aggregation is typically carried out by a control center or planning organization as represented in Figure 1. Such coordinating entities are typically found in multi-stakeholder networked systems such as Independent System Operator (ISO) in power systems and control towers in the supply chain. The aggregated dataset is then used to generate high-fidelity scenarios that can be used to solve an SO problem in a centralized fashion at the control center level. The solutions to these SO problems represent data-driven planning decisions that can be in turn used to determine an operational schedule for each stakeholder subject to constraints like reliability, safety and demand satisfaction. However, the entire data-driven planning mechanism is challenged in cases where data silos are present across various stakeholders.
@@
Therefore, the goal of SplitVAEs is to enable planning organization and control centers to generate generate stochastic decisions without incurring the burden of data processing and aggregation from multiple system stakeholders. Such planning organizations typically have regulatory oversight on the operational aspects of these systems necessitating SO problems to be solved in a high-frequency fashion. Thus, our proposed SplitVAEs framework relies on distributed memory-based aggregation primitives to conduct training without the need to move local data. As a result, the SplitVAEs algorithmic framework, by its very design is fully capable of functioning in the presence of geographically distributed, decentralized data silos as is typically found in multi-stakeholder infrastructure systems. Therefore, the SplitVAEs framework can potentially be extended to operate in hybrid edge-cloud environments based on containerized applications enabling significant reductions in data transmission. Additionally, it is capable of succinctly capturing spatiotemporal interdependencies in a decentralized manner despite the inherent non-linearities in the underlying data among multiple stakeholder subsystems while maintaining high scalability. SplitVAEs complement existing stochastic optimization frameworks by supplying high-fidelity scenarios generated from siloed datasets. More importantly, SplitVAEs are not meant to be a replacement of optimization solution methodologies. As a result, we can concisely summarize the contributions of our study as follows:
  • We design a computationally efficient learning paradigm that combines edge-based autoencoders with a server-driven variational autoencoder to jointly learn spatiotemporal interdependencies of siloed data.
@@
The remaining parts of this study are organized in the following manner. Section II provides a literature review of prior work. Sections III and IV describe the architecture and the algorithmic design of our proposed framework, respectively. Section V provides a summary of benchmark methods as well as statistical evaluation metrics. Section VI analyze the performance of the SplitVAEs framework across different application domains and diverse computational settings. Section VII serves as the concluding section of this study.
@@ II Related Works
There have been several scenario generation mechanisms that have been proposed including auto-regressive moving average (ARIMA) methods [9, 1], Gaussian copula-based approaches [10, 11] and deep generative adversarial networks (GANs) [12, 7]. A majority of these studies, however, are centralized in nature leading to computational and data privacy related concerns [7]. From a computational standpoint, centralized methods such as ARIMA and copula-based approaches rely on the construction and factorization of covariance matrices [13, 11] aided by the accumulation of stakeholder data. On the other hand, centralized GAN-based methods can be notoriously difficult to train owing to high instability and challenges in convergence [14]. As a result, scalability remains an issue with these methods, especially for large-scale networks that consist of many sources of uncertainty.
@@
Centralized methods also depend on nonlinear modeling of relationships between predictor feature variables and the generated scenarios are connected in space and time. The authors in [13] used wind speed data to construct scenarios of wind power generation via the power curve of wind turbines, while the work in [15] leveraged the nonlinear relationship between temperature and power consumption using epi-spline functions to generate useful scenarios. These approaches might become untenable, especially when integrating numerous heterogeneous sources of data, including but not limited to wind, PV farms, and alternative fuel sources such as hydrogen.
@@
On the other hand, VAE [16] serve as a potential alternative to methods that rely on non-linear modeling requirements. VAEs can help synthesize new, previously unseen scenarios by capturing the distribution of more concise, low-dimensional embeddings in a latent space. By sampling from the learned latent space distributions, VAEs can be used to generate scenarios that are spatially and temporally interdependent. Unlike GANs however, the VAEs provide a scalable, and decomposable implementation paradigm which is suited for data silos with heterogenous features as well.
@@
Different application areas have their own challenges with respect to siloed data. For example, visibility across stakeholder subsystems is a major hindrance in supply chain and logistics owing to heterogeneity and incompatibility of local data storage systems [17]. Similarly, in the medical domain, data privacy is the primary reason for the existence of data silos [18]. Privacy obstacles to scenario generation could also stem from the sensitivity of underlying datasets from the commercial standpoint [19]. Such data sharing challenges are especially prevalent in the case of variable renewable energy (VRE) stakeholders [20, 8]. Therefore, conventional methods of scenario generation face several limitations in terms of their applicability to heterogeneous data from diverse stakeholders.
@@
Split learning [21, 22, 23] presents a more generalized approach to handling siloed data sources in a decentralized, privacy-friendly fashion. Instead of relying on horizontal partitioning of data, split learning relies on model partitioning schemes that yield low-dimensional embeddings from heterogeneous data. Split learning paradigms leverage these embeddings to learn the interdependencies of the features contained in the siloed heterogeneous datasets without the need to move local data. As a result, split learning methods can potentially help resolve the data heterogeneity and data silo challenges inherent in scenario generation methods. Therefore, our approach combines the split learning paradigm with the VAE to yield a computationally efficient, privacy-preserving framework capable of delivering high-quality scenarios from decentralized and siloed datasets.
@@ III SplitVAEs Framework Design
We consider the SplitVAEs framework from the perspective of edge and server-level agents. The edge can be used to represent heterogeneous system stakeholders that govern individual data silos comprising time-series datasets. In our framework, the server houses the VAE and represents the control center or planning organization, while the edge entities consist of an autoencoder (AE) that is jointly trained in conjunction with the VAE at the server level. We begin our discussion by describing the conventional, centralized VAE (Central-VAE) formulation.
@@ III-A Centralized variational autoencoder (VAE) formulation
The Central-VAE model consumes raw data $y\in\mathbb{R}^{d}$ to yield latent space embeddings, denoted by $z\in\mathbb{R}^{s}$, with the help of a probabilistic encoder $q^{cvae}_{\phi}(z|y)$. Specifically, $q^{cvae}_{\phi}(z|y)$ initially predicts the mean and variance $(\mu_{s},\sigma_{s})$ of the latent space distribution as represented in Equation (1). The $(\mu_{s},\sigma_{s})$ are used to generate latent space representations $z$ with the help of sampling from the standard normal distribution as represented in Equation (2). The integration of a random sample $\epsilon$ is referred to as the reparametrization trick and is necessary to avoid computationally expensive simulation-driven latent space estimation. Similarly, a probabilistic decoder $p^{cvae}_{\theta}(y|z)$ is applied so as to obtain a reconstruction $\tilde{y}\in\mathbb{R}^{d}$ of the raw data $y$ represented by Equation (3).
$(\mu,\sigma)\leftarrow q^{cvae}_{\phi}(x)$ (1)
$z\leftarrow\mu+\sigma\odot\epsilon_{i},\text{ where }\epsilon_{i}\sim\mathcal{N}(0,I)$ (2)
$\tilde{x}_{i}\leftarrow p^{cvae}_{\theta}(z_{i})$ (3)
Equations (1) to (3) are considered to be one forward pass of the VAE model. On the other hand, the backward pass of the VAE is carried out with the help of a total loss function represented by Equation (4). The total loss comprises of two errors: the reconstruction error, denoted as $L^{recon}(\theta,x,z)$ and indicated by Equation (5); the Kullback-Leibler divergence, denoted as $L^{KL}(\phi,\theta,x,z)$ and derived in (6), which measures the similarity of latent space distributions based on the probabilistic encoders and decoders.
$L(\theta,\phi,x,z)=L^{recon}(\theta,x,z)+L^{KL}(\phi,\theta,x,z)\text{\quad}$ (4)
$L^{recon}(\theta,x,z)=-\mathbb{E}_{z\sim q^{cvae}_{\phi}(z|x)}\Big{(}\log p^{cvae}_{\theta}(x|z)\Big{)}\text{ }$ (5)
$L^{KL}(\phi,\theta,x,z)=D_{KL}\Big{(}q^{cvae}_{\phi}(z|x)||p^{cvae}_{\theta}(z)\Big{)}\text{\quad\quad}$ (6)
Therefore, the implementation of the Central-VAE requires the transfer of edge-level stakeholder datasets to compute the reconstruction loss as well as generate latent space embeddings, which might be infeasible due to the reasons outlined in Section I. Therefore, to learn the spatiotemporal interdependencies despite the decentralized, siloed nature of stakeholder data, the SplitVAEs framework incorporates a novel split training mechanism that leverages an edge-based AE component and a server-based VAE component.
@@ III-B Edge-level autoencoder (AE) formulation
Given the aforementioned reasons, at the edge-level, we implement an autoencoder (AE) framework comprising deep neural network (DNN) encoder and decoder models. We denote the siloed time-series data located at each edge location $n\in\mathcal{N}$ as $y^{n}$. The functional forms of the encoder and decoder components are denoted by $x^{n}\leftarrow F_{\theta^{n}_{E}}(y^{n})$ and $\tilde{y}^{n}\leftarrow G_{\theta^{n}_{D}}(x^{n})$, respectively. For the edge-based AE framework, let $x^{n},\tilde{y}^{n},\theta^{n}_{E},\theta^{n}_{D}$ be the low-dimensional embedding, the reconstructed vector, and the model parameters of the encoder and decoder, respectively. Unlike conventional AE training, a unique aspect of the SplitVAEs training mechanism is the independent training of the encoder and decoder components to holistically capture spatial and temporal interdependencies among all the edge devices.
@@ III-C Server-level VAE formulation
At the server level, we implement a conventional VAE framework. However, unlike the Central-VAE, in each training epoch, the server model consumes the set of low dimensional embeddings, denoted as $\{x^{\{1\}},\ldots,x^{\{N\}}\}$, supplied by the edge-based AEs at the various stakeholders. Let $q_{\phi}(z|x)$ be the probabilistic encoder, responsible for deriving the latent vector $z$ from the input embeddings $x$ in terms of empirical mean $\mu$ and standard deviation $\sigma$, respectively. Additionally, we denote $z\leftarrow R_{\mu,\sigma}(\epsilon)$ as the reparametrization function akin to Equation (2). The second core component is the latent vector $z$ representing the compressed data in a lower dimension. The size of $z$ becomes a crucial hyperparameter that necessitates careful tuning, in conjunction with the overall architecture of the entire network. The final component is the probabilistic decoder, denoted as $p_{\theta}(x|z)$, responsible for reconstructing a data point $x$ from a given latent vector $z$.

### v2

@@ Abstract
Stochastic optimization problems in large-scale multi-stakeholder networked systems (e.g., power grids and supply chains) rely on data-driven scenarios to encapsulate complex spatiotemporal interdependencies. However, centralized aggregation of stakeholder data is challenging due to the existence of data silos resulting from computational and logistical bottlenecks. In this paper, we present SplitVAE, a decentralized scenario generation framework that leverages variational autoencoders to generate high-quality scenarios without moving stakeholder data. With the help of experiments on distributed memory systems, we demonstrate the broad applicability of SplitVAEs in a variety of domain areas that are dominated by a large number of stakeholders. Furthermore, these experiments indicate that SplitVAEs can learn spatial and temporal interdependencies in large-scale networks to generate scenarios that match the joint historical distribution of stakeholder data in a decentralized manner. Lastly, the experiments demonstrate that SplitVAEs deliver robust performance compared to centralized, state-of-the-art benchmark methods while significantly reducing data transmission costs, leading to a scalable, privacy-enhancing alternative to scenario generation.
@@ I Introduction
footnotetext: This work was supported in part by the National Science Foundation’s SaTC program under grant CNS-2348411. The computing for this project was performed at the High Performance Computing Center at Oklahoma State University supported in part through the National Science Foundation grant OAC-1531128. (Corresponding author: H M Mohaimanul Islam).
Stochastic optimization (SO) problems are essential components in large-scale, data-driven planning and decision frameworks for multi-stakeholder-led networked systems including but not limited to energy [1], supply chain [2], spare parts management [3], healthcare [4] and logistics [5]. Good quality solutions to SO problems fundamentally require high-fidelity data-driven scenarios that effectively encapsulate spatial and temporal interdependencies from myriad sources [6]. Consequently, the generation of high-fidelity scenarios is heavily contingent on the availability of high-quality datasets that capture these interdependencies from heterogeneous data sources. Conventional scenario generation methods necessitate centralized data aggregation to capture interdependencies to drive stochastic planning decisions. These centralized aggregation schemes, however, face several computational, and logistical challenges [7, 8] due to the presence of data silos across multiple stakeholders of a networked system. Therefore, in this study, we present SplitVAE, a novel decentralized scenario generation framework leverages split-learning to train variational autoencoders (VAEs) to eliminate the demand for centralized data aggregation. SplitVAEs are compatible with pre-existing data silos, ensure scalable generation of high-fidelity scenarios and significantly reduce data transfer costs.
Conventionally, SO solution frameworks rely on the acquisition of local spatiotemporally correlated datasets from individual stakeholders. The mechanism of data aggregation is typically carried out by a control center or planning organization as represented in Figure 1. Such coordinating entities are typically found in multi-stakeholder networked systems, for instance, Independent System Operators (ISOs) in power systems and control towers in the supply chain. The aggregated dataset is then used to generate high-fidelity scenarios that can be used to solve an SO problem in a centralized fashion at the control center level. The solutions to these SO problems represent data-driven planning decisions that can be in turn used to determine an operational schedule for each stakeholder subject to constraints like reliability, safety and demand satisfaction. However, the entire data driven planning mechanism is challenged in cases where data silos are present across various stakeholders.
@@
Therefore, the goal of SplitVAEs is to enable planning organizations and control centers to generate generate stochastic decisions without incurring the burden of data processing and aggregation from multiple system stakeholders. Such planning organizations typically have regulatory oversight on the operational aspects of these systems necessitating SO problems to be solved in a high-frequency fashion. Thus, our proposed SplitVAE framework relies on distributed memory based aggregation primitives to conduct training without the need to move local data. As a result, SplitVAEs by their very design are fully capable of functioning in the presence of geographically distributed, decentralized data silos as is typically found in multi-stakeholder infrastructure systems.
As a result, SplitVAEs can potentially be extended to operate in hybrid edge-cloud environments based on containerized applications enabling significant reductions in data transmission. Additionally, it is capable of succinctly capturing spatiotemporal interdependencies in a decentralized manner despite the inherent non-linearities in the underlying data among multiple stakeholder subsystems while maintaining high scalability. SplitVAEs complement existing stochastic optimization frameworks by supplying high-fidelity scenarios generated from siloed datasets. SplitVAEs are not meant to be a replacement of stochastic optimization solution methodologies. Instead, they help generate spatiotemporally interdependent scenarios from siloed or geographically dispersed datasets, which can then be used as input for solving stochastic optimization problems. As a result, we can concisely summarize the contributions of our study as follows:
  • We design a computationally efficient learning paradigm that combines edge-based autoencoders with a server driven variational autoencoder to jointly learn spatiotemporal interdependencies of siloed data.
@@
The remaining parts of this study are organized in the following manner. Section II provides a literature review of prior work. Sections III and IV describe the architecture and the algorithmic design of our proposed framework, respectively. Section V provides a summary of benchmark methods as well as statistical evaluation metrics. Section VI analyzes the performance of the SplitVAEs framework across different application domains and diverse computational settings. Section VII serves as the concluding section of this study.
@@ II Related Works
There have been several scenario generation mechanisms that have been proposed including auto-regressive integrated moving average (ARIMA) methods [9, 10], Gaussian copula-based approaches [11, 12] and deep generative adversarial networks (GANs) [13, 7]. A majority of these studies, however, dealt with data centralization leading to computational and data privacy concerns [7]. From a computational standpoint, centralized methods such as ARIMA and copula-based approaches rely on the construction and factorization of covariance matrices [14, 12] aided by the accumulation of stakeholder data. Meanwhile, centralized GAN-based methods suffer from high instability and challenges in convergence [15] during model training. As a result, scalability remains an issue with these methods, especially for large-scale networks and systems consisting of myriad sources of uncertainty.
@@
Further, centralized methods depend on nonlinear modeling of relationships between predictor feature variables and the generated scenarios that are connected in space and time. The authors in [14] used wind speed data to construct scenarios of wind power generation via the power curve of wind turbines. The work in [16] leveraged the nonlinear relationship between temperature and power consumption using epi-spline functions to generate scenarios. These approaches assume homogenous sources of data. When integrating heterogeneous sources of data such as wind, PV farms, and alternative fuel sources like hydrogen, these approaches might become infeasible.
@@
On the other hand, VAEs [17] serve as a potential alternative to methods that rely on non-linear modeling requirements. VAEs can help synthesize new, previously unseen scenarios by capturing the distribution of more concise, low-dimensional embeddings in a latent space. By sampling from the learned latent space distributions, VAEs can be used to generate scenarios that are spatially and temporally interdependent. Unlike GANs however, the VAEs provide a scalable, and decomposable implementation paradigm which is suited for data silos with heterogenous features as well.
@@
Additionally, challenges with respect to siloed data arise in several application domains. For example, visibility across stakeholder subsystems is a major hindrance in supply chain and logistics owing to heterogeneity and incompatibility of local data storage systems [18]. Similarly, in the medical field, data privacy is the primary reason for the existence of data silos [19]. Privacy obstacles to scenario generation could potentially stem from the sensitivity of underlying datasets from the commercial standpoint [20]. Such data sharing challenges are especially prevalent in the case of variable renewable energy (VRE) stakeholders [21, 8]. Overall, conventional methods of scenario generation face several limitations in terms of their applicability to heterogeneous data and requirements for privacy preservation from diverse stakeholders.
@@
Split learning [22, 23, 24] presents a more generalized approach to handling siloed data sources in a decentralized, privacy-friendly fashion. Instead of relying on horizontal partitioning of data, split learning relies on model partitioning schemes that yield low-dimensional embeddings from heterogeneous data. Split learning paradigms leverage these embeddings to learn the interdependencies of the features contained in the siloed heterogeneous datasets without the need to move local data. As a result, split learning methods can potentially help resolve the data heterogeneity and data silo challenges that are inherent in scenario generation methods. Thus, our approach combines the split learning paradigm with the VAE to yield a computationally efficient, privacy-enhancing framework capable of delivering high-quality scenarios from decentralized and siloed datasets.
@@ III SplitVAEs Framework Design
We consider the SplitVAEs framework from the perspective of edge and server-level agents. The edge level represents heterogeneous system stakeholders that govern individual data silos comprising time series datasets. In our framework, the server entity houses the VAE and represents a planning organization, coordination agency, or a control center, while the edge entities consist of an autoencoder (AE) that is jointly trained in conjunction with the VAE at the server level. We begin our discussion by describing the conventional, centralized VAE (Central-VAE) formulation.
@@ III-A Centralized variational autoencoder (VAE) formulation
The Central-VAE model consumes data provided by edge-level AEs, denoted by $x\in\mathbb{R}^{d}$, to yield latent space embeddings, denoted by $z\in\mathbb{R}^{s}$, with the help of a probabilistic encoder, denoted by $q^{\text{cvae}}_{\phi}(z|x)$. Specifically, $q^{\text{cvae}}_{\phi}(z|x)$ initially predicts the mean and variance of the latent space distribution, denoted by $(\mu_{s},\sigma_{s})$, as represented in Equation (1). The mean and variance pair $(\mu_{s},\sigma_{s})$ are then utilized to generate latent space representations $z$ with the help of sampling from the standard normal distribution as represented in Equation (2). The integration of a random sample, denoted by $\epsilon$, is referred to as the reparametrization trick and is necessary to avoid computationally expensive simulation-driven latent space estimation [17]. Similarly, a probabilistic decoder, denoted by $p^{\text{cvae}}_{\theta}(x|z)$, is applied so as to obtain a reconstruction, $\tilde{x}\in\mathbb{R}^{d}$, of the provided data $x$ represented by Equation (3).
$(\mu_{s},\sigma_{s})\leftarrow q^{\text{cvae}}_{\phi}(x);\quad x\in\mathbb{R}^{d}$ (1)
$z\leftarrow\mu_{s}+\sigma_{s}\odot\epsilon,\text{ where }\epsilon\sim\mathcal{N}(0,I_{s})$ (2)
$\tilde{x}\leftarrow p^{\text{cvae}}_{\theta}(z);\quad z\in\mathbb{R}^{s},\tilde{x}\in\mathbb{R}^{d}$ (3)
Equations (1) to (3) are considered to be one forward pass of the VAE model. On the other hand, the backward pass of the VAE is carried out with the help of a total loss function represented by Equation (4). The total loss comprises of two errors: the reconstruction error, denoted by $L^{\text{recon}}(\theta,x,z)$ and indicated by Equation (5); the Kullback-Leibler divergence, denoted by $L^{\text{KL}}(\phi,\theta,x,z)$ and derived in (6), which measures the similarity of latent space distributions based on the probabilistic encoders and decoders.
$L(\theta,\phi,x,z)=L^{\text{recon}}(\theta,x,z)+L^{\text{\text{KL}}}(\phi,\theta,x,z)\text{\quad}$ (4)
$L^{\text{recon}}(\theta,x,z)=-\mathbb{E}_{z\sim q^{\text{cvae}}_{\phi}(z|x)}\Big{(}\log p^{\text{cvae}}_{\theta}(x|z)\Big{)}\text{ }$ (5)
$L^{\text{\text{KL}}}(\phi,\theta,x,z)=D_{\text{KL}}\Big{(}q^{\text{cvae}}_{\phi}(z|x)\big{|}\big{|}p^{\text{cvae}}_{\theta}(z)\Big{)}\text{\quad\quad}$ (6)
Therefore, the implementation of the Central-VAE requires the transfer of edge-level stakeholder datasets to compute the reconstruction loss as well as generate latent space embeddings, which might be infeasible due to the reasons outlined in Section I. Therefore, to learn the spatiotemporal interdependencies given the decentralized, siloed nature of stakeholder data, the SplitVAEs framework incorporates a novel split training mechanism that leverages an edge-based AE component and a server-based VAE component.
@@ III-B Edge-level autoencoder (AE) formulation
Given the aforementioned reasons, at the edge-level, we implement an AE-based framework comprising deep neural network (DNN) encoder and decoder models. We denote the siloed time series data located at each edge location $n\in\mathcal{N}$ as $y^{n}$. The functional forms of the encoder and decoder components are denoted by $x^{n}\leftarrow F_{\theta^{n}_{E}}(y^{n})$ and $\tilde{y}^{n}\leftarrow G_{\theta^{n}_{D}}(x^{n})$, respectively. For the edge-based AE framework, let $x^{n},\tilde{y}^{n},\theta^{n}_{E},\theta^{n}_{D}$ be the low-dimensional embedding, the reconstructed vector, and the model parameters of the encoder and decoder, respectively. Unlike conventional AE training, a unique aspect of the SplitVAEs training mechanism is the independent training of the encoder and decoder components to holistically capture spatial and temporal interdependencies among all the edge devices.
@@ III-C Server-level VAE formulation
At the server level, we implement a conventional VAE framework. However, unlike the Central-VAE, in each training epoch, the server model consumes the set of low dimensional embeddings, denoted by $\{x^{1},\ldots,x^{N-1},x^{N}\}$, supplied by the edge-based AEs at the various stakeholders. Let $q_{\phi}(z|x)$ be the probabilistic encoder, responsible for deriving the latent vector $z$ from the input embeddings $\{x\}$ in terms of empirical mean and standard deviation - denoted by $\hat{\mu}$ and $\hat{\sigma}$, respectively. Additionally, we denote $z\leftarrow R_{\hat{\mu},\hat{\sigma}}(\epsilon)$ as the reparametrization function akin to Equation (2). The second core component is the latent vector $z$ representing the compressed data in a lower dimension. The size of $z$ becomes a crucial hyperparameter that necessitates careful tuning, in conjunction with the overall architecture of the entire network. The final component is the probabilistic decoder, denoted by $p_{\theta}(x|z)$, responsible for reconstructing a data point $x$ from a given latent vector $z$.